Parth (parthkohli):

Factor \(\large x^{2n} + x^{n + 2}\) I get \(\large x^{n}(x^{n} + x^{2})\) Correct, right?

OpenStudy (anonymous):

yeah u r right..

Parth (parthkohli):

Thanks

OpenStudy (unklerhaukus):

but\[ x^{n}(x^{n} + x^{2})=x^{n}x^n+x^nx^2=x^{n^2}+x^{n+2}\]

Parth (parthkohli):

How did you get the n^2? \(x^n \times x^n = x^{n + n} = x^{2n}\)

OpenStudy (anonymous):

unkle is wrong

OpenStudy (anonymous):

it cannnt be n^2.... its simply 2n

OpenStudy (unklerhaukus):

oh yeah, my mistake